a
34. (III) An air bubble at the bottom of a lake 41.0 m deep has a volume of 1.00 cm. If the temperature at the bottom is 5.5°C and at the top 18.5°C, what is the radius of the bubble just before it reaches the surface?
(I) If the air pressure at a particular place in the mountains is 0.80 atm, estimate the temperature at which water boils.
56. (II) If the air pressure at a particular place in the mountains s 0.80 atm, estimate the temperature at which water boils.
Answer #1

34. (III) Pressure at the bottom of a lake which will be given as -

P1 = Patm + Pw

P1 = Patm + \rhow g h

where, \rhow = density of water = 1000 kg/m3

g = acceleration due to gravity = 9.8 m/s2

h = depth of a lake = 41 m

Patm = atmospheric pressure = 1.013 x 105 Pa

then, we get

P1 = [(1.013 x 105 Pa) + (1000 kg/m3) (9.8 m/s2) (41 m)]

P1 = 5.031 x 105 Pa

Using a combined gas law & we get

P1 V1 / T1 = P2 V2 / T2

V2 = P1 V1 T2 / P2 T1

where, P2 = pressure at the surface of a lake = 1.013 x 105 Pa

V1 = volume of an air bubble = 1 cm3

T1 = temperature at the bottom = 5 0C = 278.65 K

T2 = temperature at the top = 18.5 0C = 291.65 K

then, we get

V2 = [(5.031 x 105 Pa) (1 cm3) (291.65 K)] / [(1.013 x 105 Pa) (278.65 K)]

V2 = 5.198 cm3

Radius of the bubble just before it reaches the surface which will be given by -

V2 = (4/3) \pi r23

(5.198 cm3) = (4/3) (3.14) r23

r23 = [(5.198 cm3) / (4.186)]

r23 = 1.24176 cm3

r2 = 1.07 cm

56. (II) Using a Gay-lussac's law & we get

P1 / T1 = P2 / T2

where, P1 = air pressure = 0.80 atm

P2 = atmospheric pressure = 1 atm

T1 = temperature at which water boils = ?

T2 = boiling temperature of water = 100 0C = 373 K

then, we get

T1 = [(0.80 atm) (373 K)] / (1 atm)

T1 = 298.4 K